OpenStudy (anonymous):

how would you classify the following reaction that forms hydrogen and oxygen gases \[2H _{2}O(l)\rightarrow2H _{2(g)+O _{2(g)}}\] a. double displacement b. synthesis c. decomposition d. combustion

OpenStudy (anonymous):

b

OpenStudy (anonymous):

It is definitely not synthesis. Let me give you some generic examples and you can determine the answer from those. A, B, X and Y will be individual species. a) \[XY + AB = XB + AY\] b) \[X + Y = XY\] c)\[XY = X + Y\] d) \[XY + A = AX + Y\]